I have no idea how to figure out what the equations are for this question:
The longest cable suspension bridge in the world is the Akashi Kaikyo Bridge in Japan. It has one long span and two short spans. Each short span is half the length of the long span. The total length of the bridge is 3560m. What is the length of each span?
i know that one equation should be x+y= 3560
and the other equation should be something along the lines of x+2y but i'm not sure what it should equal and how to make sure it represent half the length.
could someone explain this to me?

Answer by checkley77(12569)

Let x be the shorter spans.
Let y be the longer span.
x=y/2
2x+y=3,560 Noqw replace x in this equation with (y/2).
2(y/2)+y=3,560
2y/2+y=3,560
y+y=3,560
2y=3.560
y=3,560/2
y=1,780 ans. for the longer span.
x=1,780/2
x=890 ans. for the 2 shorter spans.
--------------------------------------
x/2+x/2+x=3560
2x=3560
x=3560/2
x=1780 m is the length of the long spans.
1780/2=890 m is the length of the 2 shorter spans.
Proof:
890+890+1780=3560
3560=3560
----------------------------------
I always try to take the shortest method to get the solution.